OpenStudy (anonymous):

graph the line through (5.-3) with m=2

OpenStudy (ranga):

First locate the point (5,-3) on the graph paper. slope m = 2 slope = rise / run. so slope = 2 means for every 1 step you take in the x-direction, the y will rise by +2. If you take 4 steps in the x-direction then y must increase by +8. So from the point (5,-3) take a few steps along the x axis and take twice as many steps up along the y-axis to get another point. Draw a traight line passing through those two points.

OpenStudy (anonymous):

the y intercept form is -13

OpenStudy (anonymous):

y=mx+b

OpenStudy (anonymous):

line equations have a point-slope form\[y-y_0=m(x-x_0)\]

OpenStudy (anonymous):

That might come in handy

OpenStudy (anonymous):

so y=2x-13 how to graph this??

OpenStudy (anonymous):

slope is rise/run, so with m=2 you go up 2 units for every 1 horizontal unit

OpenStudy (anonymous):

-13 is your y-intercept

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

solution y=mx+b -3=2(5)+b -3=10+b b=-3-10 b=-13 y=2x-13

OpenStudy (anonymous):



OpenStudy (anonymous):

my solution is correct or wrong??

OpenStudy (anonymous):

i started on the y-intercept, and just drew a line with slope 2\[\frac{rise}{run}=\frac{2}{1}\]

OpenStudy (anonymous):

oh i'll check it gimme a sec

OpenStudy (anonymous):

yeah it's correct

OpenStudy (ranga):

Your straight line equation is correct. But here they want you to graph it.

OpenStudy (anonymous):

yes